Fermi National Laboratory
 

Accelerator Update for June 18, 2007 - June 20, 2007

 

Monday June 18

The day shift began with Operations monitoring Tevatron (TeV) store 5500 (40.12E30) and the Recycler stash (148.2E10), with the Antiproton Source (Pbar) stacking (55.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:28 AM, a Pbar Accumulator RF station (ARF1) tripped off. It reset.

At 9:33 AM, BRF11 tripped off. It reset.

At 10:12 AM, the MI beam permit tripped due to a pulse of beam in a bad position. The permit reset.

At 11:21 AM, a MI kicker (I:KPS1A) tripped off. It reset.

The Neutron Therapy Facility (NTF) reported at 12:04 PM that they had completed treating their last patient.

Operations terminated store 5500 at 12:29 PM.

Operations put the TeV into shot setup at 12:45 PM.

At 2:55 PM, BRF17 tripped off on vacuum pressure high indication coincidentally with a nearby lightning strike. It reset.

MI experts began a MIRF2 study at 3:18 PM.

Operations established store 5501 at 3:20 PM with an initial luminosity of 166.37E30.

At 3:22 PM, MIRF13 tripped off. It reset.

At 3:52 PM, the TeV quenched at sector C1. Store 5501 was lost. Operators investigated.

Operations reported at 4 PM that the Recycler had stashed antiproton twice during the day shift.

Operations reported at 4:30 PM that the TeV sector C1 cryo system had resumed normal operations. Cryo system technicians reported that they had found a number of refrigeration building roof hatches open; they were open due to the heat. The cryo electronics at C1 had gotten soaked during a thunderstorm and that was the cause of the quench.

At 6:31 PM, the MI beam permit tripped off due to a pulse of beam in a bad position. It reset.

Operations put the TeV through a dry squeeze at 7:03 PM.

At 7:17 PM, a TeV sector D3 vacuum pump (D3RP7) tripped off. It reset, but then tripped off again. After consulting with an expert, operators managed to reset the pump at 7:39 PM. It tripped again at 8:27 PM, but reset okay. Operators contacted an expert.

TeV experts began a study period at 8:28 PM.

At 8:34 PM, Operations reported that nearly 50% of the MI beam was falling out. Operators contacted an expert who investigated.

At 10:12 PM, Operations reported that Linac was having a vernier regulation problem. Operators investigated, but the problem cleared itself before they discovered the cause.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Tuesday June 19

The midnight shift began with Operations monitoring the stash (161.2E10), with Pbar stacking (24.6E10), and with NuMI and MiniBooNE taking beam.

Operations put the TeV into shot setup at 12:18 AM.

At 1:34 AM, Operations extracted the second transfer of antiprotons out of the Recycler to send to the TeV, but it fell out of the MI. The rest of the transfers were okay.

At 2:09 AM, the MI beam aborted on a negative reading from a beam monitor. Operators investigated and found that the humidity in the MI electronics room had spiked at the same time and that was the probable cause.

At 2:13 AM, Operations reported that there were MI abort kicker problems. Operators investigated and reset the kicker (I:KPS4B).

Operations reported that the humidity in the MI electronics room had come down by 2:31 AM.

Operations established store 5503 at 3:02 AM with an initial luminosity of 107.78E30.

At 3:12 AM, I:KPS1A tripped off. It reset.

At 3:50 AM, the MI beam permit dropped due to a KPS4B “trigger not ready” indication. Operators reset the permit.

At 4:12 AM, LRF4 tripped off. It reset.

At 4:29 AM, the Pbar beam permit tripped off due to a low reading on HV102R. It reset.

Operations reported at 7:13 AM that operators had searched and secured the Meson enclosures M01 through M05 and MT6.

Operations reported at 8:01 AM that the Recycler had stashed two times during the midnight shift.

At 8:53 AM, a TeV sector B11 separator sparked and caused a 30% loss of store 5503’s luminosity.

Operations terminated store 5503 at 11:30 AM.

Operations put the TeV into shot setup at 11:43 AM.

Operations established store 5504 at 1:51 PM with an initial luminosity of 138.6E30.

At 3:17 PM, Operations reported a problem with the TeV sector F4 Dewar. The Cryo Coordinator disabled the Dewar controls and sent cryo system techs to investigate.

Operations reported at 3:59 PM that the Recycler had stashed three times during the day shift.

Operations established beam to the Meson M01 absorber at 5:57 PM. Operations ran beam for five minutes and then shut down the beam line.

Operations reported at midnight that the Recycler had stashed four times during the evening shift.

Wednesday June 20

The midnight shift began with Operations monitoring store 5504 (49.06E30) and the stash (185.7E10), with Pbar stacking (14.2E10), and with NuMI and MiniBooNE taking beam.

Operations reported at 8:00 AM that the Recycler had stashed three times during the midnight shift.

Operations terminated store 5504 at 8:01 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/21/2007    email Fermilab